U have 40 tickets 30 are $150 8 are 300 how much r the last 2 so the average is $200 for all the tickets
Answer by hamsanash1981@gmail.com(151)

average of 30 ticket is 150, then total = 30*150 = 4500
similarly,
8* 300 = 2400, average of 40 = 200, total = 8000 last two tickets 2x
4500 + 2400 + 2x = 8000
2x = 8000- 6900
2x= 1100
x = 550

the last 2 tickets are 550